# Context-Selective State Space Models: Feedback is All You Need

### v1

@@ 6 Conclusion
In this work we propose a new context-selective state-space model (COFFEE) in which a state-feedback mechanism is designed to obtain a significant advantage in the model expressivity with a low number of trainable parameters while remaining suitable for parallel implementation. In the tests conducted, the COFFEE model largely outperforms S6 in both the induction head and the MNIST when the two are used within the same simple architecture. So far, COFFEE is still a proof of principle, since its performance in more complex scenarios (as, for example, the long-range arena) has not yet been tested. Of course, these scenarios require a Mamba-like architecture with several COFFEE modules. Nonetheless, the accuracy that is achieved with a single COFFEE module and a limited set of parameters is remarkable.
We are currently working on incorporating the COFFEE module in a suitable architecture, and compare it with Mamba and other, non state-space models on more challenging and diverse benchmarks. The numerical feasibility and efficiency of such implementation is granted by the chosen structure of COFFEE: since ensures a diagonal Jacobian, it remains amenable to parallelization (using the techniques recently proposed in [33]).
@@ C.1 Simplified Dynamics via Linearization
involves $A^{(i)}$, making it inherently feature-specific. This motivates assigning each subsystem its own time-invariant matrices $B^{(i)}$ and $C^{(i)}$. We therefore equip each subsystem with feature-specific $B^{(i)}$ and $C^{(i)}$, both time-invariant. [^5] Time variation arises solely through $\Delta_{i}(k)$. This yields:
@@
This matches S6, which has $nD$ parameters in each of $W_{B}$ and $W_{C}$, $nD$ for $\{A^{(i)}\}$, and $D^{2}$ for $W_{D}$, again totaling $3nD+D^{2}$. We remark that these counts refer only to the SSM block. The full architecture (presented in Appendix E) includes an additional $|M|D$ parameters from the embedding matrix, for a total of $3nD+D^{2}+|M|D$, where $M$ is the vocabulary size.
@@ C.3 Eliminating Redundancy via Change of Basis
Appendix D Induction Heads Task
@@ D.1 Task description
Appendix E Core Module Architecture
@@
    The core module receives a sequence or a batch of sequences as input. Each sequence is built according to the rules presented in Appendix D. The input can be regarded as an array of dimension [$B$, $L$], where $B$ denotes the batch size and $L$ the length of each sequence within the batch.  [^6] More generally, we use the notation $d_{1}$,…, $d_{l}$] to indicate an array having $l$ dimensions, where the first dimension is $d_{1}$ and the $l$-th dimension is $d_{l}$
@@
    $\textrm{emb}():M\rightarrow\mathbb{R}^{D}$
@@
    Notice that, the domain of the $\textrm{emb}()$ map is $M$. Since we assume that $V\subseteq M$ ($V$ is the vocabulary used to build sequences), this transformation is well-defined: it guarantees that every symbol in any input sequence has a corresponding embedding vector.
@@
Appendix F Towards a mechanistic interpretation of COFFEE: auxiliary material
F.1 Simplified Version of the Induction Head Task: IH0
@@
An embedding in $\mathbb{R}^{2}$ allows us to employ only two one-dimensional state-space models. It is hence possible to visualize the trajectories of both states on a Cartesian plane. The choice of setting $L_{\text{seq}}=4$ and using a vocabulary consisting of only three symbols, significantly reduced the number of possible sequences that could be generated according to the rules described in Appendix D.
@@
F.2 COFFEE simplified model and state evolution for IH0
@@
Appendix G Results on the Induction Head Task (IH)
@@
G.1 Experiment settings
We do not rely on a predefined dataset with a fixed set of sequences. For instance, by setting the sequence length to $L_{\text{seq}}=16$ and defining the vocabulary as $V=\{1,2,3,4,5,6,7\}$, the total number of possible sequences that can be generated, according to the rules described in Appendix D, exceeds several hundred billion.
@@ G.2 Models Initialization
Appendix H MNIST Architecture and Results
@@ H.2 Architecture Description
H.2.1 Limitations of the sMNIST Task
@@
H.2.2 Our Architecture
Our model is designed as a general tool for sequence processing. In general, when dealing with sequences, we take a sequence of symbols from some vocabulary, associate an embedding vector to each symbol, and then process the sequence of embedding vectors with the model. For instance, this is done for the induction head task presented in Appendix D.
@@ H.2.2 Our Architecture
With these parameters, we obtain a loss of $0.107$ and a accuracy of $96.6\%$.
@@
and we obtained a loss of $1.891$ with a accuracy of $28.2\%$.
@@
In this case, the architecture has $10^{\prime}085$ parameters, and we obtained a loss of $1.876$ with a accuracy of $28.6\%$.

### v2

@@ 7 Conclusion
In this work, we propose a new context-selective state-space model (COFFEE) in which a nonlinear state-feedback mechanism is designed to enhance model expressivity while remaining parameter-efficient and parallelizable. In the considered benchmarks (induction head and MNIST), COFFEE consistently outperforms S6 when the two models are instantiated within the same architectural setting. Ablation studies clearly identify the introduction of nonlinear feedback as the key source of such an advantage.
At this stage, COFFEE is still a proof of principle, since its performance in more complex benchmarks, such as the Long Range Arena, has not yet been tested. Extending COFFEE to these settings will naturally require a Mamba-like architecture composed of multiple COFFEE modules. Nonetheless, the strong accuracy achieved by a single COFFEE module with few learnable parameters provides encouraging evidence for the scalability of the proposed approach.
Building on these results, ongoing work is focused on integrating COFFEE into deeper architectures and conducting systematic comparisons with Mamba and other non–state-space models on more challenging and diverse benchmarks. Crucially, the numerical feasibility and computational efficiency of these extensions are supported by the structural properties of COFFEE: the resulting diagonal Jacobian ensures that training remains amenable to parallelization.
@@ C.1 Simplified Dynamics via Linearization
involves $A^{(i)}$, making it inherently feature-specific. This motivates assigning each subsystem its own time-invariant matrix $B^{(i)}$. We additionally assign feature-specific $C^{(i)}$ to maintain symmetry in the model. We therefore equip each subsystem with feature-specific $B^{(i)}$ and $C^{(i)}$, both time-invariant. [^7] Time variation arises solely through $\Delta_{i}(k)$. This yields:
@@
This matches S6, which has $nD$ parameters in each of $W_{B}$ and $W_{C}$, $nD$ for $\{A^{(i)}\}$, and $D^{2}$ for $W_{D}$, again totaling $3nD+D^{2}$. We remark that these counts refer only to the SSM block. The full architecture (presented in Appendix F) includes an additional $|M|D$ parameters from the embedding matrix, for a total of $3nD+D^{2}+|M|D$, where $M$ is the vocabulary size.
@@ C.3 Eliminating Redundancy via Change of Basis
Appendix D Parallelization
In this section, we provide details on the parallelization technique adopted to efficiently parallelize our model. The following exposition is based on [27] and [13].
[…]
Our goal is to find the unique trajectory $x^{*}(1),\dots,x^{*}(L)$ such that:
$\begin{cases}x^{*}(1)=f(x(0))\\ x^{*}(2)=f(x^{*}(1))\\ \vdots\\ x^{*}(L)=f(x^{*}(L-1))\end{cases}$
If we define the so-called residual as:
$r(x_{1:L})\;:=\;[x(1)-f(x(0)),x(2)-f(x(1)),...,x(L)-f(x(L-1))]\in\mathbb{R}^{L\tilde{n}}$
mathematically, we can define a vector-valued function $r:\mathbb{R}^{L\tilde{n}}\rightarrow\mathbb{R}^{L\tilde{n}}$, and our objective is to find the zero of this function. One approach to solve this problem is to apply Newton’s method.
We begin by selecting an initial guess $x_{1:L}^{(0)}$, representing a tentative trajectory of the nonlinear state-space model at the zeroth iteration. Then, we iteratively update this guess until convergence by performing the following steps:
  1. 1.
    Given the (i1)(i-1)-th solution $x^{(i-1)}_{1:L}$ perform a first order Taylor expansion of $r(\cdot)$ centered at $x^{(i-1)}_{1:L}$
  2. 2.
    From the previous point we obtain $r(x_{1:L})=r(x^{(i-1)}_{1:L})+J_{r}(x^{(i-1)}_{1:L})(x_{1:L}-x^{(i-1)}_{1:L})+o(\cdot)$
  3. 3.
    Solve $0=r(x^{(i-1)}_{1:L})+J_{r}(x^{(i-1)}_{1:L})(x_{1:L}-x^{(i-1)}_{1:L})$ in $x_{1:L}$
  4. 4.
    Set the solution to previous equation as $x^{(i)}_{1:L}$
Concretely, each iteration involves solving the following linear system:
$J_{r}(x^{(i-1)}_{1:L})(x_{1:L}-x^{(i-1)}_{1:L})=-r(x^{(i-1)}_{1:L})$
By evaluating $J_{r}(x^{(i-1)}_{1:L})$ we obtain:
$\begin{bmatrix}I_{n}&0&0&\cdots&0&0\\ -\dfrac{\partial f}{\partial x}\!\bigl(x^{(i-1)}(1)\bigr)&I_{n}&0&\cdots&0&0\\ 0&-\dfrac{\partial f}{\partial x}\!\bigl(x^{(i-1)}(2)\bigr)&I_{n}&\cdots&0&0\\ \vdots&\vdots&\ddots&\ddots&\vdots&\vdots\\ 0&0&\cdots&-\dfrac{\partial f}{\partial x}\!\bigl(x^{(i-1)}(L-2)\bigr)&I_{n}&0\\ 0&0&\cdots&0&-\dfrac{\partial f}{\partial x}\!\bigl(x^{(i-1)}(L-1)\bigr)&I_{n}\end{bmatrix}$
which is a block bidiagonal matrix in $\mathbb{R}^{L\tilde{n}\times L\tilde{n}}$. Although this matrix is invertible, directly solving the linear system by computing $J_{r}(x^{(i-1)}_{1:L})^{-1}$ is computationally intractable at large scale.
However, we can exploit the block bidiagonal structure to efficiently solve the system. In particular, by defining
$\Delta x^{(i)}(k)\;:=\;x^{(i)}(k)-x^{(i-1)}(k)$
and expanding the equations accordingly, we can avoid the full matrix inversion while still performing the Newton update. In fact, by forward substitution on:
$J_{r}(x^{(i-1)}_{1:L})(x_{1:L}-x^{(i-1)}_{1:L})=-r(x^{(i-1)}_{1:L})$
we obtain the following first order linear recurrence:
$\begin{cases}\Delta x^{(i)}(1)=-r_{1}(x^{(i-1)}_{1:L})\\ \Delta x^{(i)}(k)=\dfrac{\partial f}{\partial x}(x^{(i-1)}(k-1))\Delta x^{(i)}(k-1)-r_{k}(x^{(i-1)}_{1:L})\end{cases}$
where
$r_{k}(x^{(i-1)}_{1:L})\;:=\;x^{(i-1)}(k)-f(x^{(i-1)}(k-1)),\quad k=1,\dots,L.$
To solve this first-order linear recurrence, we employ Blelloch’s parallel scan algorithm [5].
According to [5], we need to store the pairs
$e_{k}\;:=\;\Bigg(\frac{\partial f}{\partial x}(x^{(i-1)}(k)),\;-r_{k}(x^{(i-1)}_{1:L})\Bigg)$.
Since $\frac{\partial f}{\partial x}(x^{(i-1)}(k))$ is generally a dense matrix in $\mathbb{R}^{\tilde{n}\times\tilde{n}}$ and there are $L$ such matrices to store, the memory requirement scales as $\mathcal{O}(L\tilde{n}^{2})$. Furthermore, combining the various $e_{k}$ according to Blelloch’s algorithm requires $L$ matrix multiplications of dense $\tilde{n}\times\tilde{n}$ matrices, resulting in a total computational cost of $\mathcal{O}(L\tilde{n}^{3})$.
To mitigate these computational and memory demands, [13] proposed a modification in which only the diagonals of the Jacobians are considered. In this case, $\frac{\partial f}{\partial x}(x^{(i-1)}(k))$ becomes a diagonal matrix in $\mathbb{R}^{\tilde{n}\times\tilde{n}}$. Consequently, storing the pairs
$e_{k}\;:=\;\Bigg(\frac{\partial f}{\partial x}(x^{(i-1)}(k)),\;-r_{k}(x^{(i-1)}_{1:L})\Bigg)$
[…]
This modified algorithm is referred to as quasi-DEER in [13].
Another contribution of [13] concerns the improved numerical stability of the algorithm. The key idea is to reformulate the trajectory evaluation problem as an optimization problem. By doing so, it becomes evident that the previously described solution corresponds to a Gauss-Newton method. Introducing a trust region for the update steps transforms this into the so-called Levenberg-Marquardt algorithm.
The optimization problem obtained by applying the Levenberg-Marquardt algorithm is convex and satisfies Slater’s constraint qualification, which guarantees strong duality. The approach then consists of treating the unique Lagrange multiplier as a hyperparameter of the model and minimizing only the associated Lagrangian. When minimizing the Lagrangian with fixed Lagrange multipliers, one can observe, as noted in [38], a connection to the MAP estimate of $x_{1:L}$ in a linear Gaussian state-space model. This estimate can be computed via a Kalman smoother.
In this context, the Kalman smoother can be parallelized as described in [37]. Specifically, by considering a set $\mathcal{F}$ consisting of appropriate conditional densities and likelihoods, and by defining a suitable associative operator $\otimes$, it can be shown that $(\mathcal{F},\otimes)$ forms a monoid. This structure allows, once again, to leverage Blelloch’s parallel scan algorithm to evaluate the Kalman smoother in $\mathcal{O}(\log L)$ steps on a parallel machine.
In [13], this algorithm is referred to as Evaluating Levenberg-Marquardt with Kalman (ELK).
Appendix E Induction Head Task
@@ E.1 Task description
Appendix F Core Module Architecture
@@
    The core module receives a sequence or a batch of sequences as input. Each sequence is built according to the rules presented in Appendix E. The input can be regarded as an array of dimension [$B$, $L$], where $B$ denotes the batch size and $L$ the length of each sequence within the batch.  [^9] More generally, we use the notation $d_{1}$,…, $d_{l}$] to indicate an array having $l$ dimensions, where the first dimension is $d_{1}$ and the $l$-th dimension is $d_{l}$
@@
    $\textrm{emb}:M\rightarrow\mathbb{R}^{D}$
@@
    Notice that, the domain of the emb map is $M$. Since we assume that $V\subseteq M$ ($V$ is the vocabulary used to build sequences), this transformation is well-defined: it guarantees that every symbol in any input sequence has a corresponding embedding vector.
@@
Appendix G Towards a mechanistic interpretation of COFFEE: auxiliary material
G.1 Simplified Version of the Induction Head Task: IH0
@@
An embedding in $\mathbb{R}^{2}$ allows us to employ only two one-dimensional state-space models. It is hence possible to visualize the trajectories of both states on a Cartesian plane. The choice of setting $L_{\text{seq}}=4$ and using a vocabulary consisting of only three symbols, significantly reduced the number of possible sequences that could be generated according to the rules described in Appendix E.
@@
G.2 COFFEE simplified model and state evolution for IH0
@@
Appendix H Results on the Induction Head Task (IH)
@@
H.1 Experiment settings
We do not rely on a predefined dataset with a fixed set of sequences. For instance, by setting the sequence length to $L_{\text{seq}}=16$ and defining the vocabulary as $V=\{1,2,3,4,5,6,7\}$, the total number of possible sequences that can be generated, according to the rules described in Appendix E, exceeds several hundred billion.
@@ H.2 Models Initialization
Appendix I MNIST Architecture and Results
@@ I.2 Architecture Description
I.2.1 Limitations of the sMNIST Task
@@
I.2.2 Our Architecture
Our model is designed as a general tool for sequence processing. In general, when dealing with sequences, we take a sequence of symbols from some vocabulary, associate an embedding vector to each symbol, and then process the sequence of embedding vectors with the model. For instance, this is done for the induction head task presented in Appendix E.
@@ I.2.2 Our Architecture
With these parameters, we obtain a loss of $0.107$ and a accuracy of $96.6\%$ with standard deviation of 0.1 (repeated ten times).
@@
and we obtained a loss of $1.891$ with a accuracy of $28.2\%$ with standard deviation of 0.1 (repeated ten times).
@@
In this case, the architecture has $10^{\prime}085$ parameters, and we obtained a loss of $1.876$ with a accuracy of $28.6\%$ with standard deviation of 0.1 (repeated ten times).